OpenStudy (anonymous):

there is 2 spinners. one containing numbers 1,2,3,4,5 and second spinner containing colors: red, red, green, blue what is the probability of spinning an even number and the color red? i know there is 2 reds and 2 even numbers so i thought the answer would be 4/9 but its not one of the answers in the choices i have.. possible answers: 1/10 1/5 4/5 3/20

OpenStudy (kropot72):

There are two even numbers in the set of five numbers. So the probability of an even number is 2/5. The probability of red is 2/4. Therefore the probability of spinning an even number and the color red is given by: \[\frac{2}{5} \times \frac{2}{4}=\frac{1}{5}\]